OpenStudy (anonymous):

Raul has to choose 3 of the 5 toppings for his frozen yogurt. How many different combinations of toppings are possible? A. 10 B. 20 C. 60 D. 120

OpenStudy (anonymous):

This is just 5 choose 3

OpenStudy (anonymous):

Do you know how to calculate choose?

OpenStudy (anonymous):

No..lol

OpenStudy (anonymous):

n choose k = n! / ( k! (n-k)! )

OpenStudy (anonymous):

So 5 choose 3 = 5! / (3! 2!)

OpenStudy (anonymous):

That is confusing, I don't understand.

OpenStudy (anonymous):

Do you know what 5! is?

OpenStudy (anonymous):

No. I don't know how to do that, is it 5x4x3x2x1?

OpenStudy (anonymous):

They properly didn't prepare you for this question.

OpenStudy (anonymous):

Yes, that is correct 5! = 5x4x3x2x1

OpenStudy (anonymous):

Okay so I just plug that into the whole equation?

OpenStudy (anonymous):

So now you need to do 3! and 2!

OpenStudy (anonymous):

Then add them all together?

OpenStudy (anonymous):

then plug them into 5! / (3!2!)

OpenStudy (anonymous):

I'm still confused...

OpenStudy (anonymous):

Okay what is 5!?

OpenStudy (anonymous):

I have already did that, the equation doesn't make sense.

OpenStudy (anonymous):

okay it is 5! / (3! x 2!)

OpenStudy (anonymous):

Multiply bottom ones, then divide the result from 5!

OpenStudy (anonymous):

Is it 60? I'm on a timed practice.

OpenStudy (anonymous):

no. 5! = 120, 3! = 6, 2! = 2 5 choose 3 = 120 / (6 * 2) = 120 / 12 = 10

OpenStudy (anonymous):

Thank you.